OpenStudy (anonymous):

use Newton Method to approximate (25)^(1/8) upto 6 decimal places.

OpenStudy (mimi_x3):

x = (25)^(1/8) x^8 = 25 x^8 - 25 = 0

OpenStudy (mimi_x3):

f(x) = x^8 - 25 f'(x) = 8x^7

OpenStudy (mimi_x3):

then use newtons method of approx.

OpenStudy (mimi_x3):

and try x_1 = 1

OpenStudy (anonymous):

thanks :)

OpenStudy (mimi_x3):

np (: